Two point charges are placed as follows: charge q1 = -1.50 nC is at y = +6.00 m and charge 

q2 = +3.20 nC is at the origin. What is the total force (magnitude and direction) exerted by 

these two charges on a negative point charge q3 = -5.00 nC located at (2.00 m, -4.00 m)?


1
Expert's answer
2022-03-23T07:15:44-0400

The net force


"\vec F_{net}=k\frac{q_1q_3}{r_{13}^3}\vec r_{13}+k\frac{q_2q_3}{r_{23}^3}\vec r_{23}""\vec r_{13}=(2,-10),\quad r_{13}=\sqrt{2^2+(-10)^2}=\sqrt{104},\\
\vec r_{23}=(2,-4),\quad r_{23}=\sqrt{2^2+(-4)^2}=\sqrt{20}.""\vec F_{net}=9*10^9*\frac{(-1.5*10^{-9})*(-5.0*10^{-9})}{(\sqrt{104})^3}*(2,-10)\\
+9*10^9*\frac{(3.20*10^{-9})*(-5.0*10^{-9})}{(\sqrt{20})^3}*(2,-4)""\vec F_{net}=6.36*10^{-11}*(2,-10)\\
-1.61*10^{-9}*(2,-4)=(-3.1*10^{-9},5.8*10^{-9})"

The magnitude of force


"F_{net}=\sqrt{(-3.1)^2+5.8^2}*10^{-9}=6.6*10^{-9}\:\rm N"

Direction


"\theta=\tan^{-1}\frac{F_y}{F_x}=\tan^{-1}\frac{5.8}{-3.1}=118^\circ\: \rm from\; +x"